How do you write the equation of the indicated conic in standard form?
Ellipse: Vertices:(-4,6),(8,6) Co-vertices:(2,8),(2,4)

Answer by lwsshak3(11628)

How do you write the equation of the indicated conic in standard form?
Ellipse: Vertices:(-4,6),(8,6) Co-vertices:(2,8),(2,4)
**
This is an ellipse with horizontal major axis:
Its equation: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k)=(x,y) coordinates of center
y-coordinate of center=6
x-coordinate of center=2 (8-4)/2=2 (midpoint formula)
center: (2,6)
length of horizontal major axis=12 (-4 to 8)=2a
a=6
a^2=36
..
length of co-vertices of minor axis=4 (4 to 8)=2b
b=2
b^2=4
..
Equation:
(x-2)^2/36+(y-6)^2/4=1